the sum of the interior angles of a polygon is 540 degrees how many sides does it have
Answer by ikleyn(52792)

.

The sum of interior angles of a n-gon is  180*(n-2) degrees.


In your case, equation for unknown number of sides (of vertices) is

    180*(n-2) = 540  degrees.


Divide both sides by 180

    n-2 = 540/180 = 3

     n            = 3 + 2.


ANSWER.  n = 5:  your polygon is a pentagon.

Solved.